### v1

@@ 5 Results
Tables 2 and 3 present the model performance for predicting fire occurrence (FO) and burned area (BA) across four forecast lead times: $1,7,15$, and $31$ days. For each target, we report two complementary metrics: a global metric (x) computed over all departments combined, and an area-normalized score (y) computed as the area under the performance curve per department. The global metrics reflect aggregate accuracy on the entire test set, whereas the area-normalized metrics treat each department equally. This separation highlights spatial generalization by evaluating how well each model predicts fires in every department, in addition to its overall performance.
- Model Performance on global metrics
The CatBoost ensemble emerges as the top performer in the fire risk prediction benchmark, achieving the highest scores on both targets. For example, at the longest forecast horizon (31 days), CatBoost reaches an F1 score of 0.71 for fire occurrence—the best among all models—along with an intersection-over-union (IoU) of 0.45 for occurrence and 0.48 for burned area. Most other methods attain slightly lower but comparable global scores: XGBoost and even logistic regression perform nearly as well (with 31-day F1 scores around the high 0.60s), whereas deep learning models generally lag behind these simpler approaches.Longer prediction lead times consistently yield better model performance.
- Model Performance on Area-Normalized Across Departments
FWI yields higher area-normalized F1 scores across departments than the learned models. For instance, at a 31-day horizon, FWI attains an area-normalized F1 of 0.55, exceeding CatBoost’s 0.53. However, its IoU is comparatively low. This discrepancy likely stems from FWI’s lack of calibration to actual fire counts or regional conditions, which causes it to broadly overpredict risk. While such coverage yields high recall, FWI’s low overall accuracy (e.g., 0.15 global F1 at 1 day) limits its practical utility. In contrast, deep learning models often outperform tree-based models on area-normalized F1 and IoU, indicating better geographic generalization. Although area scores often correlate with recall, we observe marginal cases (e.g., GRU vs. DilatedCNN at 7 days, where both achieved comparable area-normalized scores—0.38 and 0.37—but different recall: 0.75 and 0.71, respectively), suggesting that recall alone does not fully explain generalization performance.
- Binary vs. Multi-classification
Binarizing the fire prediction task reveals clear trade-offs, with binary models achieving higher area-normalized metrics but lower global metrics such as Catboost at 15 days with 0.64, 0.5 in binary versus 0.68, 0.43 for the multi-class version—reflecting improved regional coverage. An exception exists at 15 days, with the MLP binary obtaining lower scores in both metrics. Figure 5 shows that for GRU, the multi-class model excels in high-fire departments but underperforms beyond 25 fires, where binary performance remains more stable. Similar curves for CatBoost and MLP are available in the supplementary materials. Figure 6 illustrates Catboost prediction in Bouches-du-Rhône and Charente. Multi-class predictions in Bouches-du-Rhône capture fire severity, while in Charente, binary models yield better detection despite missing fire count granularity. These results confirm that multi-class models perform best in high-activity zones, whereas binary models offer more consistent coverage in lower-risk areas. It seems that this is due to classes 2, 3, and 4 are difficult to predict due to their rarity, where classification, by assigning everything to class 1, somewhat mitigates this issue.
- Fire Occurrence vs. Burned Area
Predictions for BA consistently show better area-normalized performance than those for FO. For instance, at a 7-day forecast horizon, models achieve an IoU of approximately 0.42 for BA, compared to about 0.39 for FO, indicating higher spatial accuracy in predicting burned areas. This difference in predictability is largely attributed to class distribution: the positive class 1 (minor fires) occurs much more frequently in BA than in FO, reducing both class imbalance and prediction variance. Furthermore, the clustering approach aggregates extreme fire sizes into a single highest category. As a result, most BA predictions fall within class 0 (no fire) and class 1, leading to a narrower and more stable range of outputs. This reduced output space simplifies the prediction task, making BA significantly easier to predict than FO.
- Combining all samples’ performance and departments’ performance
The final model performance was evaluated by multiplying global accuracy (x) with area-normalized performance (y), resulting in a composite score that captures both overall effectiveness and spatial consistency. The final score is illustrated in Figure 7. For IoU on FO, GRU achieves the highest performance at 0 days, CatBoost leads at 7 and 31 days, and MLP performs best at 15 days with CatBoost being second. Models are followed by GraphCast (0 days), DilatedCNN (7 days), and LSTM (31 days). Regarding the F1 score for FO, GraphCast performs best at 0 days, while CatBoost Binary leads at 7 and 15 days. However, as the prediction horizon increases, the probability of at least one fire occurring becomes less interpretable. The highest score at 31 days is achieved by the multi-class version of CatBoost. For IoU on Burned Area (BA), GRUGNN performs best at 0 days (with Logistic Regression as runner-up), while CatBoost leads at all longer horizons, with GNN-based models (GRUGNN, GraphCast) and Logistic Regression consistently ranking among the top performers. These trends indicate that deep learning models are more effective for short-term forecasts, whereas CatBoost provides stable and high accuracy across all time horizons. When comparing GRUGNN to GRU, we found that the GNN-based model tends to perform better at longer time horizons (15–31 days) for FO, and at 31 days for BA. For short-term predictions, the GRU model remains superior. As discussed further in Section 6, this may suggest that decoupling spatial from temporal features helps the model better capture spatial patterns when they are most significant. That said, the performance differences remain relatively small, and this observation should therefore be considered preliminary.
@@
Table 2: Model comparison for Fire Occurrence (FO) and Burned Area (BA) at 0-day and 7-days prediction horizons. “LR” stands for Logistic Regression. We do not show FWI performance on BA as it is made for fire occurrence. In addition, IoU is not shown for binary models as it is not comparable with multi-classification. The first value is the global score, and the second value (when noted) is the area-normalized score.
Model FO (0 Day) BA IoU
F1 Rec Prec IoU
FWI 0.15, 0.14 0.84 0.09 0.05, 0.05 -
LR 0.42, 0.14 0.50 0.36 0.24, 0.08 0.23, 0.08
XGBoost 0.42, 0.14 0.52 0.36 0.24, 0.08 0.24, 0.07
CatBoost 0.43, 0.13 0.52 0.37 0.24, 0.07 0.23, 0.06
MLP 0.42, 0.14 0.60 0.32 0.24, 0.08 0.23, 0.06
GRUGNN 0.41, 0.14 0.66 0.24 0.24, 0.08 0.24, 0.08
GraphCast 0.41, 0.16 0.61 0.31 0.23, 0.09 0.22, 0.07
LSTM 0.41, 0.15 0.64 0.30 0.23, 0.09 0.23, 0.07
GRU 0.43, 0.15 0.57 0.35 0.24, 0.09 0.23, 0.07
DilatedCNN 0.40, 0.16 0.54 0.32 0.22, 0.09 0.22, 0.07
CatBoost Binary 0.42, 0.15 0.63 0.31 - -
MLP Binary 0.40, 0.15 0.58 0.32 - -
GRU Binary 0.37, 0.17 0.76 0.24 - -
Model FO (7 Days) BA IoU
F1 Rec Prec IoU
FWI 0.38, 0.38 0.84 0.29 0.17, 0.17 -
LR 0.63, 0.33 0.68 0.59 0.39, 0.20 0.41, 0.20
XGBoost 0.63, 0.34 0.68 0.59 0.39, 0.20 0.42, 0.20
CatBoost 0.64, 0.35 0.69 0.59 0.40, 0.21 0.42, 0.20
MLP 0.61, 0.37 0.73 0.52 0.34, 0.21 0.41, 0.19
GRUGNN 0.59, 0.37 0.80 0.47 0.34, 0.21 0.40, 0.21
GraphCast 0.61, 0.34 0.66 0.56 0.37, 0.20 0.40, 0.20
LSTM 0.58, 0.36 0.72 0.47 0.37, 0.22 0.40, 0.19
GRU 0.60, 0.37 0.75 0.50 0.37, 0.22 0.41, 0.21
DilatedCNN 0.60, 0.37 0.71 0.52 0.38, 0.22 0.39, 0.20
CatBoost Binary 0.60, 0.40 0.69 0.59 - -
MLP Binary 0.54, 0.39 0.86 0.40 - -
GRU Binary 0.58, 0.38 0.75 0.48 - -
Table 3: Model comparison for Fire Occurrence (FO) and Burned Area (BA) at 15-days and 31-days prediction horizons.
Model FO (15 Days) BA IoU
F1 Rec Prec IoU
FWI 0.47, 0.47 0.81 0.39 0.22, 0.22 -
LR 0.67, 0.44 0.72 0.62 0.41, 0.27 0.44, 0.27
XGBoost 0.67, 0.42 0.71 0.64 0.41, 0.25 0.44, 0.26
CatBoost 0.68, 0.43 0.73 0.67 0.43, 0.26 0.45, 0.28
MLP 0.63, 0.47 0.82 0.51 0.40, 0.29 0.45, 0.25
GRUGNN 0.64, 0.45 0.80 0.53 0.40, 0.27 0.42, 0.28
GraphCast 0.63, 0.46 0.80 0.52 0.39, 0.28 0.43, 0.29
LSTM 0.63, 0.44 0.76 0.54 0.40, 0.27 0.43, 0.26
GRU 0.64, 0.45 0.76 0.56 0.40, 0.27 0.44, 0.27
DilatedCNN 0.64, 0.44 0.72 0.57 0.40, 0.26 0.42, 0.26
CatBoost Binary 0.64, 0.50 0.86 0.51 - -
MLP Binary 0.63, 0.46 0.78 0.53 - -
GRU Binary 0.61, 0.46 0.79 0.50 - -
Model FO (31 Days) BA IoU
F1 Rec Prec IoU
FWI 0.47, 0.55 0.81 0.47 0.25, 0.25 -
LR 0.69, 0.54 0.77 0.64 0.43, 0.33 0.46, 0.35
XGBoost 0.71, 0.50 0.71 0.71 0.44, 0.31 0.46, 0.33
CatBoost 0.71, 0.53 0.76 0.67 0.45, 0.34 0.48, 0.35
MLP 0.69, 0.53 0.77 0.63 0.42, 0.33 0.46, 0.34
GRUGNN 0.68, 0.53 0.81 0.59 0.43, 0.33 0.45, 0.35
GraphCast 0.67, 0.47 0.68 0.67 0.41, 0.29 0.45, 0.35
LSTM 0.68, 0.54 0.77 0.61 0.43, 0.34 0.44, 0.35
GRU 0.67, 0.52 0.76 0.61 0.43, 0.33 0.44, 0.34
DilatedCNN 0.66, 0.53 0.72 0.61 0.41, 0.33 0.41, 0.34
CatBoost Binary 0.65, 0.58 0.91 0.51 - -
MLP Binary 0.66, 0.57 0.84 0.55 - -
GRU Binary 0.67, 0.57 0.82 0.57 - -
Figure 5: Area‑normalized F1 comparison for the GRU model: multi‑class versus binary formulations as the number of departments varies.
Figure 6: Catboost prediction summary at 0 days horizons for Bouches du Rhone in multi-classification (a) and binary (b), and Charente in multi-classification (c) and binary (d).
Figure 7: Final score achieved by each model.
@@ 6 Features importance
We investigated feature importance by computing SHAP values on the multi-class CatBoost model. Figure 8 shows the top 15 features for fire occurrence across different forecast horizons (0, 7, 15, and 31 days). A corresponding figure for burned area prediction is provided in the supplementary materials.
At 0 day, fire prediction is primarily driven by history and short-term signals. Encoded temporal variables, spatial clustering, and immediate weather conditions dominate the model’s decision-making. Fire danger indices, relative humidity, and vegetation-specific factors—such as the presence of pine trees—contribute meaningfully, highlighting the importance of real-time environmental conditions. By 7 days, we found similar results, but Historical features become more influential, along with temporal indicators such as the month. By 15 days, the shift toward structural and seasonal features continues. Fire history, time-of-year encodings, and forest topographic indicators become central to the model’s predictions. By 31 days, compared to shorter-term forecasts, the model relies almost exclusively on historical, static, or slowly evolving features. Vegetation and NDVI gain importance, and socio-cultural variables also start to contribute.
Figure 8: Top 15 features computed on multi-classification Catboost model on different time horizons for fire occurrence. ID encoder corresponds to Department ID
7 Future work
In this section, we outline several directions for future work aimed at improving current model performance.
@@ 7 Future work
Current risk classes are based solely on historical fire occurrences at the departmental level. Incorporating seasonal data and clustering similar regions may help capture risk dynamics that go beyond purely historical patterns.
@@
Our analysis of feature importance revealed that historical variables are strong predictors in regions with frequent past fire activity. While these features help the model capture seasonal fire trends, their static nature may bias predictions toward past patterns and overlook emerging fire-prone areas. Reducing their influence during training could help the model prioritize more transferable features. We also acknowledge that 2D convolutional networks such as ResNet and ConvLSTM were not included in our benchmark. Although effective for spatial tasks like fine-scale map prediction, their high computational cost made them unsuitable for our department-scale feature set. While they seem irrelevant at short-term forecast, spatial features for long-term forecasts gain in influence as highlighted in Section 6. We plan to explore 2D CNNs (after dimensionality reduction) in future work.
@@ 8 Conclusion
We present the first nationwide benchmark for wildfire risk prediction across all metropolitan departments in France. By analyzing this approach, we report multiple findings. A five-level ordinal classification task provides richer and more actionable information than a binary approach, particularly for high risk departments. However, multi-class models are sensitive to the clustering scheme: K-means clustering for burned area tends to compress moderate risk levels while exaggerating extreme events, making this approach currently less suitable than regression. Global versus local evaluation reveals that models may appear accurate in aggregate while underperforming in sparsely burned departments, highlighting the importance of combining overall and area-normalized metrics. CatBoost consistently performs best overall, while deep learning models tend to outperform in short-term (day-ahead) forecasts. Decoupling temporal and spatial features appears promising for enhancing spatial interpretation but requires further investigation. Feature importance analysis reveals an over-reliance on historical fire signals, which limits the model’s ability to generalize to emerging risk areas. Future work will explore tail-robust discretization strategies, ordinal-aware loss functions, federated learning, and 2D CNN architectures following dimensionality reduction, to capture effectively fine-scale spatial patterns.

### v2

@@ 4 Results
Tables 3 present the models performance for predicting FO and BA. Table 4 present the models performance for predicting fire risk in each department. This is give by equation 1 for T being each studied area (here French department), and score a score function.
$Score\;=\;\frac{\displaystyle\int_{0}^{T}\mathrm{score}(t)\,\mathrm{d}t}{\displaystyle\int_{0}^{T}\mbox{max\_score}\,\mathrm{d}t}$ (1)
This score analyze how well the model is able to predict everywhere. The idea behind this equation is to compare the area under the curve representing each department’s score, normalized by the maximum score (here 1 for each metric). By selecting the departments where at least one fire occurred in 2023, we can evaluate the model’s ability to make predictions within each department. Note that this equation works when the score has a maximum value and this value is the best value possible. For that reason we don’t compute the area score for auoc (minimum is better).
4.1 Model Performance on Global Metrics
Across all models and tasks, recall values are consistently higher than precision, both in global and area-normalized evaluations. For instance, in FO prediction, CatBoost reaches a recall of 0.51 versus a precision of 0.37, while GRU attains 0.56 recall for only 0.31 precision. This behavior is desirable, specially in low risk region where missing an actual fire is more critical than producing false alarms. We observe that FWI achieves the highest recall with the lowest AUOC, while the other metrics remain very low. This suggests that the metric, in its original implementation, has rather limited relevance for predicting occurrences
4.2 Model-Specific Observations
Among all tested models, CatBoost and GraphCastGRU achieve the best F1 scores (both 0.43). GraphCastGRU offers superior recall (0.60 vs. 0.33 precision), a higher IoU (0.25 compared to CatBoost’s 0.24), and better spatial generalization. From the area-normalized evaluation, GraphCastGRU maintains F1 = 0.16, IoU = 0.09 for FO, and IoU = 0.08 for BA.
4.3 Comparison of FO and BA
When comparing the average IoU values of Fire Occurrence (FO) and Burned Area (BA), FO appears slightly easier to predict. Global IoU scores for FO range between 0.22–0.24 across models, compared to 0.21–0.23 for BA. This indicates that FO, as a target, is more predictable, whereas BA involves higher uncertainty linked to the dynamics of fire spread. There are several reasons that may explain this phenomenon: (1) Burned Area is inherently a more stochastic target compared to occurrence, as it depends not only on ignition but also on spread dynamics influenced by local conditions. (2) As observed earlier, the class distribution of Burned Area is more imbalanced than that of Fire Occurrence, leading models to underpredict higher BA classes and thus lowering performance metrics.
4.4 Binary vs. Multi-class Evaluation
Binary and multi-class models achieve relatively close F1 scores. For FO, binary CatBoost reaches F1 = 0.13 under area normalization, while multi-class CatBoost scores 0.13 as well. GRU binary and multi-class both yield F1 = 0.16. However, precision-recall trade-offs differ: binary classifiers tend to favor higher recall (e.g., 0.26 for GRU binary) at the cost of lower precision (0.13), while multi-class models remain more balanced (precision around 0.13–0.15, recall around 0.22–0.27).
This confirms that probability prediction alone is not sufficient at this scale: the ordinal structure of fire risk levels carries valuable information that can be better leveraged. Moreover, both global and area-normalized results show that further improvements require going beyond pure probability outputs.
Overall, these findings underline the necessity of multi-metric evaluation. Recall-oriented behavior is well aligned with operational needs, but only models such as GraphCastGRU combine high F1 (0.43), strong IoU (0.25), positioning it as the most promising approach for localized fire risk prediction in France. The generalization accross department is a still a main challenge. Figure 3 shows predictions obtain by GraphCastGRU in Bouches du Rhône. Based on theses predictions we can clearly see the challenging task of predicting highest classes.
@@ 4.3 Comparison of FO and BA
Table 3: Model comparison for FO and BA at 0-day “LR” stands for Logistic Regression. We do not show FWI performance on BA as it is made for FO. In addition, IoU is not shown for binary models as it is not comparable with multi-classification. Best values are shown in bold.
Model FO BA
F1 Prec Rec IoU auoc IoU auoc
FWI 0.17 0.10 0.87 0.06 0.619 - -
LR 0.39 ± 0.01 0.28 ± 0.00 0.65 ± 0.01 0.23 ± 0.00 0.66 ± 0.00 0.24 ± 0.00 0.71 ± 0.00
catboost 0.43 ± 0.00 0.37 ± 0.00 0.51 ± 0.00 0.24 ± 0.00 0.72 ± 0.00 0.25 ± 0.00 0.71 ± 0.00
xgboost 0.42 ± 0.01 0.32 ± 0.01 0.60 ± 0.01 0.24 ± 0.00 0.70 ± 0.00 0.25 ± 0.00 0.71 ± 0.01
GRU 0.40 ± 0.01 0.31 ± 0.01 0.56 ± 0.04 0.23 ± 0.00 0.70 ± 0.01 0.23 ± 0.01 0.71 ± 0.01
LSTM 0.40 ± 0.01 0.29 ± 0.01 0.61 ± 0.01 0.23 ± 0.00 0.69 ± 0.01 0.23 ± 0.01 0.70 ± 0.01
MLP 0.39 ± 0.01 0.31 ± 0.02 0.52 ± 0.02 0.22 ± 0.01 0.71 ± 0.01 0.21 ± 0.01 0.69 ± 0.00
DilatedCNN 0.35 ± 0.01 0.24 ± 0.01 0.61 ± 0.02 0.20 ± 0.01 0.70 ± 0.00 0.21 ± 0.01 0.71 ± 0.01
graphCastGRU 0.43 ± 0.01 0.33 ± 0.02 0.60 ± 0.05 0.25 ± 0.00 0.70 ± 0.01 0.25 ± 0.01 0.71 ± 0.01
catboost Binary 0.43 ± 0.00 0.37 ± 0.01 0.50 ± 0.01 - - - -
GRU Binary 0.40 ± 0.01 0.30 ± 0.01 0.62 ± 0.01 - - - -
NetMLP Binary 0.39 ± 0.01 0.30 ± 0.01 0.55 ± 0.02 - - - -
Table 4: Area model comparison for FO and BA at 0-day prediction.
Model FO BA
F1 Prec Rec IoU IoU
FWI 0.14 0.08 0.83 0.058 -
LR 0.15 ± 0.01 0.14 ± 0.01 0.29 ± 0.01 0.09 ± 0.00 0.08 ± 0.00
catboost 0.13 ± 0.00 0.13 ± 0.00 0.18 ± 0.00 0.07 ± 0.00 0.08 ± 0.00
xgboost 0.15 ± 0.00 0.13 ± 0.01 0.24 ± 0.00 0.09 ± 0.00 0.08 ± 0.01
GRU 0.16 ± 0.02 0.13 ± 0.02 0.24 ± 0.04 0.09 ± 0.00 0.08 ± 0.00
LSTM 0.16 ± 0.01 0.13 ± 0.01 0.27 ± 0.01 0.09 ± 0.00 0.09 ± 0.01
MLP 0.15 ± 0.01 0.15 ± 0.01 0.22 ± 0.02 0.08 ± 0.01 0.09 ± 0.00
DilatedCNN 0.16 ± 0.01 0.12 ± 0.01 0.31 ± 0.03 0.09 ± 0.01 0.09 ± 0.01
graphCastGRU 0.16 ± 0.00 0.14 ± 0.01 0.23 ± 0.04 0.09 ± 0.00 0.08 ± 0.01
catboost Binary 0.13 ± 0.01 0.13 ± 0.00 0.18 ± 0.00 - -
GRU Binary 0.16 ± 0.00 0.13 ± 0.01 0.26 ± 0.01 - -
NetMLP Binary 0.16 ± 0.01 0.15 ± 0.02 0.24 ± 0.01 - -
Figure 3: GraphCastGRU prediction for Bouches du Rhone in 2023.
@@ 5 Features importance
We investigated feature importance by computing SHAP values on the multi-class CatBoost model. Figure 4 shows the top 15 features for FO. A corresponding figure for BA prediction is provided in the supplementary materials.
Fire prediction is primarily driven by historical and short-term signals. Encoded temporal variables, spatial clustering, and immediate weather conditions dominate the model’s decision-making. Fire danger indices, relative humidity, and vegetation-specific factors—such as the presence of pine trees—also contribute meaningfully.
Historical features are strong predictors in historically active regions but can hurt generalization in low-risk or emerging areas. While they help the models to easily detect the seasonal trend of fire, their static nature biases the model toward past patterns, potentially missing new fire-prone zones.
Figure 4: Top 15 features computed on multi-classification Catboost model on different time horizons for FO. ID encoder corresponds to Department ID
6 Discussion
In this section, we outline several possible directions for future work aimed at improving current model performance.
@@ 6 Discussion
Current risk classes are based solely on historical FOs and BA at the departmental level. Incorporating seasonal data and clustering similar regions may help capture risk dynamics that go beyond purely historical patterns.
@@
Our analysis of feature importance revealed that historical variables are strong predictors in regions with frequent past fire activity. While these features help the model capture seasonal fire trends, their static nature may bias predictions toward past patterns and overlook emerging fire-prone areas. Reducing their influence during training could help the model prioritize more transferable features. We also acknowledge that 2D convolutional networks such as ResNet and ConvLSTM were not included in our benchmark. Although effective for spatial tasks like fine-scale map prediction, their high computational cost made them unsuitable for our department-scale feature set. We plan to explore 2D CNNs (after dimensionality reduction) in future work.
@@ 7 Conclusion
In this work, we introduced a new nationwide database for wildfire risk prediction, specifically aligned with the departmental organization of French firefighting services and enriched with local variables such as calendar features. While this dataset can serve broader research domains, we proposed the first benchmark for predicting both fire occurrence and burned area using an ordinal classification scheme, a novel approach in this context. Our experiments highlight three key findings: (1) multi-class prediction provides a clear advantage over binary formulations at this scale, (2) GraphCastGRU achieves the best overall performance across metrics which corroborate with findings of [6], and (3) the current ordinal scheme is less suitable for burned area prediction. Future directions include exploring ordinal-aware loss functions, applying federated learning to assess generalization, and testing more computationally demanding architectures such as CNNs and Transformers to better capture spatial and temporal dependencies.